standfast01:

Ginny has a sheet of plywood that is 4 feet by 8 feet. She wants to saw the plywood into pieces 1 foot high and 2 feet long for signs. How many signs can she make with the sheet of plywood?

standfast01:

@Elsa213

standfast01:

@slipknot3066

standfast01:

Okay

Elsa213:

Google says 16 butt I dunno ef es right. e.e

standfast01:

Okay hold up

standfast01:

Yeah it is right

Elsa213:

Congratulations. c'x

slipknot3066:

Leave it tew mamma Elsa tew save de day

standfast01:

Wait but how did you get that Did google explain it to you?

Elsa213:

Please dun. Lemme copy pasta du explanation

Elsa213:

4*8 = 32 1*2 = 2 32/2 = 16 Whale, now et seems logical. e.e

slipknot3066:

Oof

Elsa213:

I cri

standfast01:

Oh okay thx

JustSaiyan:

*coughs* maybe recheck.

Elsa213:

^ Math masta

JustSaiyan:

How many signs can you make out of a sheet of wood 4 by 8. You divide

JustSaiyan:

or subtract. Pick your poison

standfast01:

Subtract...

JustSaiyan:

The answer would be 4. You'd be cutting this piece of wood in equal parts. 1 feet by 2 feet goes into 4 feet by 8 feet 4 times.

standfast01:

Oh okay...